### v1

@@ 2 Literature Review
Social media bots, or fondly called “bots”, refer to social media users that are software-controlled and can automatically perform a series of tasks. These types of user accounts are of keen interest to the social cybersecurity community because they have been observed to perform malicious activities online, which can affect the peace of society. They have observed to be used to infiltrate political discourse and spread misinformation. During the 2010 US midterm elections, social bots were already observed to have been flooding the social media space with their support for some candidates and smear their opponents by injecting thousands of tweets pointing to websites with fake news ferrara2016rise . Bots working together in a coordinated fashion have also been known to apply social pressure on to other users, causing them to change their opinion towards key topics. This was observed in the case of the 2021 coronavirus vaccination debate, where users surrounded by coordinating active bots change their stance towards the vaccine, potentially resulting in an anti-vaccine stance and the real-world refusal of the vaccine ng2022pro .
@@ 3.1 Data
User name
Screen Name
Description
Posts
User Metadata
botometer-feedback-2019 yang2019arming
529 (27)
Y
Y
P
P
P
botwiki-2019 yang2019arming
704 (100)
Y
Y
P
P
P
cresci-rtbust-2019 mazza2019rtbust
759 (52)
Y
Y
P
P
P
cresci-stock-2018 cresci2018fake
25987 (71)
Y
Y
P
P
P
midterms-2018 yang2020scalable
50538 (84)
Y
Y
P
P
P
political-bots-2019 yang2019arming
62 (100)
Y
Y
Y
N
Y
reddit-2022
667 (75)
Y
N
Y
Y
Y
instagram-2022
1862 (100)
Y
Y
P
N
Y
@@ 3.2 BotBuster For Everyone: Ensemble Bot Detection
Username
string entropy, #uppercase letters, #lowercase letters, #digits, #punctuations, #emojis, #hashtags
[…]
Screenname
string entropy, #uppercase letters, #lowercase letters, #digits, #punctuations, #emojis, #hashtags, #words
[…]
Description*
TF-IDF
[…]
69.84
81.59
79.26
User Metadata
#followers, #following, #listed, #posts, #likes, protected, verified
[…]
Posts**
TF-IDF, #likes, #retweets, #replies, #quotes
56.37
[…]
79.97
[…]
@@ 3.3 Model Evaluation
botometer-feedback-2019
57.60 (61.44)
59.05 (71.07)
83.08 (100)
botwiki-2019
53.12 (90.34)
48.12 (92.90)
91.60 (100)
cresci-rtbust-2019
61.89 (74.97)
69.43 (78.78)
71.65 (100)
cresci-stock-2018
37.20 (40.57)
39.25 (47.03)
74.61 (100)
midterms-2018
13.20 (11.26)
14.15 (1.31)
85.23 (100)
political-bots-2019
0 (0)
17.33(20.60)
74.54 (100)
verified-2019
88.60 (100)
35.50 (98.15)
99.57 (100)
reddit-2022
0.30 (0)
0 (0)
35.68 (100)
instagram-2022
0 (0)
0 (0)
60.26 (100)
Average
34.62 (42.06)
31.42(45.52)
75.14 (100)
@@
BotHunter
99.15
45.98
41.48
49.02
Botometer
91.38
33.02
35.98
30.86
BotBuster For Everyone
100
57.32
44.62
50.69
@@
botometer-feedback-2019
86.55 (54.30)
83.08 (100)
botwiki-2019
95.20 (84.94)
91.60 (100)
cresci-rtbust-2019
73.67 (41.90)
71.65 (100)
cresci-stock-2018
78.54 (26.40)
74.61 (100)
midterms-2018
87.93 (14.55)
85.23 (100)
political-bots-2019
NA (0)
74.54 (100)
verified-2019
99.67 (88.45)
99.57 (100)
reddit-2022
NA (0)
35.68 (100)
instagram-2022
NA (0)
60.26 (100)
Average
86.93 (51.76)
75.14 (100)
@@ Appendix A Full Accuracy Metrics
% processed
Accuracy
Micro-F1 score
Macro-F1 score
botometer-feedback-2019
100
83.08
83.08
78.81
botwiki-2019
100
91.60
34.34
48.55
cresci-rtbust-2019
100
71.65
70.83
70.13
cresci-stock-2018
100
74.61
74.43
74.38
midterms-2018
100
85.23
89.73
49.46
political-bots-2019
100
74.54
78.64
68.45
verified-2019
100
99.57
99.42
50.29
reddit-2022
100
34.68
51.20
33.86
instagram-2022
100
60.26
69.67
37.61
Average
100
50.69
44.62
56.84
@@
% processed
Accuracy (processed)
Accuracy (overall)
Micro-F1 score (Overall)
Macro-F1 score (Overall)
botometer-feedback-2019
61.44
74.10
57.60
70.65
69.75
botwiki-2019
90.34
53.13
53.12
69.39
34.69
cresci-rtbust-2019
74.97
62.90
91.89
62.98
65.87
cresci-stock-2018
40.57
37.36
37.20
48.09
35.93
midterms-2018
100
11.26
15.30
13.20
9.02
political-bots-2019
0
0
0
0
0
verified-2019
88.60
100
100
100
100
reddit-2022
0
0
0
0
0
instagram-2022
0
0
0
0
0
Average
52.45
57.13
40.37
40.01
34.55
@@
% processed
Accuracy (processed)
Accuracy (overall)
Micro-F1 score (Overall)
Macro-F1 score (Overall)
botometer-feedback-2019
71.07
53.68
59.05
60.95
63.89
botwiki-2019
92.90
92.89
48.12
50.07
46.37
cresci-rtbust-2019
78.78
38.12
69.43
69.42
69.38
cresci-stock-2018
47.03
38.12
39.25
44.10
41.39
midterms-2018
47.03
11.90
14.15
12.40
7.22
political-bots-2019
1.31
20.60
17.33
21.48
20.96
verified-2019
20.60
30.20
35.50
25.71
27.53
reddit-2022
0
0
0
0
0
instagram-2022
0
0
0
0
0
Average
45.52
31.72
31.43
31.57
30.75

### v2

@@ 2 Literature Review
Social media bots, or fondly called “bots”, refer to social media users that are software-controlled and can automatically perform a series of tasks. These types of user accounts are of keen interest to the social cybersecurity community because they have been observed to perform malicious activities online, which can affect the peace of society. They have observed to be used to infiltrate political discourse and spread misinformation. During the 2010 US midterm elections, social bots were already observed to have been flooding the social media space with their support for some candidates and smear their opponents by injecting thousands of tweets pointing to websites with fake news ferrara2016rise . Bots are also used by countries for digital diplomacy, to put forth a desired narrative facing the online public ng2023deflating ; jacobs2023tracking . Bots working together in a coordinated fashion have also been known to apply social pressure on to other users, causing them to change their opinion towards key topics. This was observed in the case of the 2021 coronavirus vaccination debate, where users surrounded by coordinating active bots change their stance towards the vaccine, potentially resulting in an anti-vaccine stance and the real-world refusal of the vaccine ng2022pro .
@@ 3.1 Data
User name Screen Name Description Posts User Metadata
botometer-feedback-2019 yang2019arming 529 (27) Y Y P P P
botwiki-2019 yang2019arming 704 (100) Y Y P P P
cresci-rtbust-2019 mazza2019rtbust 759 (52) Y Y P P P
cresci-stock-2018 cresci2018fake 25987 (71) Y Y P P P
midterms-2018 yang2020scalable 50538 (84) Y Y P P P
political-bots-2019 yang2019arming 62 (100) Y Y Y N Y
reddit-2022 667 (75) Y N Y Y Y
instagram-2022 1862 (100) Y Y P N Y
@@ 3.2 BotBuster For Everyone: Ensemble Bot Detection
Username string entropy, #uppercase letters, #lowercase letters, #digits, #punctuations, #emojis, #hashtags 75.81R,IG 75.94IG 72.72IG 72.14R,IG
Screenname string entropy, #uppercase letters, #lowercase letters, #digits, #punctuations, #emojis, #hashtags, #words 75.69R,IG 79.54IG 72.08R,IG 71.57IG
Description* TF-IDF 70.48IG 69.84 81.59 79.26
User Metadata #followers, #following, #listed, #posts, #likes, protected, verified 100IG 74.80IG 100R,IG 100R,IG
Posts** TF-IDF, #likes, #retweets, #replies, #quotes 56.37 81.02R 79.97 79.02R
@@ 3.3 Model Evaluation
botometer-feedback-2019 57.60 (61.44) 59.05 (71.07) 83.08 (100)
botwiki-2019 53.12 (90.34) 48.12 (92.90) 91.60 (100)
cresci-rtbust-2019 61.89 (74.97) 69.43 (78.78) 71.65 (100)
cresci-stock-2018 37.20 (40.57) 39.25 (47.03) 74.61 (100)
midterms-2018 13.20 (11.26) 14.15 (1.31) 85.23 (100)
political-bots-2019 0 (0) 17.33(20.60) 74.54 (100)
verified-2019 88.60 (100) 35.50 (98.15) 99.57 (100)
reddit-2022 0.30 (0) 0 (0) 35.68 (100)
instagram-2022 0 (0) 0 (0) 60.26 (100)
Average 34.62 (42.06) 31.42(45.52) 75.14 (100)
@@
BotHunter 99.15 45.98 41.48 49.02
Botometer 91.38 33.02 35.98 30.86
BotBuster For Everyone 100 57.32 44.62 50.69
@@
botometer-feedback-2019 86.55 (54.30) 83.08 (100)
botwiki-2019 95.20 (84.94) 91.60 (100)
cresci-rtbust-2019 73.67 (41.90) 71.65 (100)
cresci-stock-2018 78.54 (26.40) 74.61 (100)
midterms-2018 87.93 (14.55) 85.23 (100)
political-bots-2019 NA (0) 74.54 (100)
verified-2019 99.67 (88.45) 99.57 (100)
reddit-2022 NA (0) 35.68 (100)
instagram-2022 NA (0) 60.26 (100)
Average 86.93 (51.76) 75.14 (100)
@@ Appendix A Full Accuracy Metrics
% processed Accuracy Micro-F1 score Macro-F1 score
botometer-feedback-2019 100 83.08 83.08 78.81
botwiki-2019 100 91.60 34.34 48.55
cresci-rtbust-2019 100 71.65 70.83 70.13
cresci-stock-2018 100 74.61 74.43 74.38
midterms-2018 100 85.23 89.73 49.46
political-bots-2019 100 74.54 78.64 68.45
verified-2019 100 99.57 99.42 50.29
reddit-2022 100 34.68 51.20 33.86
instagram-2022 100 60.26 69.67 37.61
Average 100 50.69 44.62 56.84
@@
% processed Accuracy (processed) Accuracy (overall) Micro-F1 score (Overall) Macro-F1 score (Overall)
botometer-feedback-2019 61.44 74.10 57.60 70.65 69.75
botwiki-2019 90.34 53.13 53.12 69.39 34.69
cresci-rtbust-2019 74.97 62.90 91.89 62.98 65.87
cresci-stock-2018 40.57 37.36 37.20 48.09 35.93
midterms-2018 100 11.26 15.30 13.20 9.02
political-bots-2019 0 0 0 0 0
verified-2019 88.60 100 100 100 100
reddit-2022 0 0 0 0 0
instagram-2022 0 0 0 0 0
Average 52.45 57.13 40.37 40.01 34.55
@@
% processed Accuracy (processed) Accuracy (overall) Micro-F1 score (Overall) Macro-F1 score (Overall)
botometer-feedback-2019 71.07 53.68 59.05 60.95 63.89
botwiki-2019 92.90 92.89 48.12 50.07 46.37
cresci-rtbust-2019 78.78 38.12 69.43 69.42 69.38
cresci-stock-2018 47.03 38.12 39.25 44.10 41.39
midterms-2018 47.03 11.90 14.15 12.40 7.22
political-bots-2019 1.31 20.60 17.33 21.48 20.96
verified-2019 20.60 30.20 35.50 25.71 27.53
reddit-2022 0 0 0 0 0
instagram-2022 0 0 0 0 0
Average 45.52 31.72 31.43 31.57 30.75